OpenStudy (anonymous):

Fill in the blank: 2t+8 is greater than or equal to -4(t+1) t is greater than or equal to ___

OpenStudy (anonymous):

all you need to do is solve for T. would you like me to show you how?

OpenStudy (anonymous):

Yes please

OpenStudy (yanasidlinskiy):



OpenStudy (anonymous):

t is greater than or equal to 2 becuase you first use distributive property on the right side which gives you -4t + -4 then you would subtract 4 from both sides and get 2t+12 is greater than or equal to -4t and then -2t from both sides and then just divide each side by -6 and switch the sign

OpenStudy (igreen):

\(2t + 8 \ge -4(t + 1)\) Distribute -4 into the parenthesis: \(2t + 8 \ge -4t - 4\) Subtract 8 to both sides: \(2t \ge -4t - 12\) Add 4t to both sides: \(6t \ge -12\) Now divide 6 to both sides, can you do that?

OpenStudy (igreen):

@styles0201

OpenStudy (anonymous):

You get t is grater than or equal to -2 ? @iGreen

OpenStudy (anonymous):

*greater

OpenStudy (anonymous):

2T+8(=/greater)-4(T+1) distribute first. 2T+8=-4T-4 then get the T alone by subtracting -4T from each side. 6T+8=-4 then subtract 8 from both sides. 6T=-12 last divide by 6 to find out what T is equal to or greater than. T=-2

OpenStudy (anonymous):

I was confused on whether it was 2 or -2. Thank you! @mason437